Jonh is tree times as old as sarah, last year jonh was four times as lod as sarah. So how old is sarah and jonh now??
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

Sarah be x years
John = 3x years
..
last year
sarah was x-1
John was 3x-1
..
3x-1=4(x-1)
3x-1=4x-4
add 1 to both sides
3x+1-1=4x+1-4
3x=4x-3
add -4x to both sides
3x-4x=4x-4x-3
-x=-3
divide by -1
x= 3
sarah = 3 years
John = 3x = 9 years

Answer by stanbon(75887)

John is three times as old as Sarah.
J = 3S
Last year John was four times as oid as Sarah.
(J-1) = 4(S-1)
-------------------------
So how old is Sarah and John now??
---
Substitute for "J" and solve for "S":
(3S-1) = 4(S-1)
3S-1 = 4S-4
S = 3 yrs. old (Sarah's age now)
------------------
Since J= 3S, J = 3*3 = 9 yrs old (John's age now)
======================================================
Cheers,
Stan H.
========